OpenStudy (anonymous):

2 word problems from my final review? 1. A jar contains 10 blue marbles, 4 red marbles, and 8 white marbles. What are the odds of drawing a blue marble from the bag? Answer: 5/11 2) A ladder 12 feet long makes an angle of 67 degrees with the ground as it leans against a store. How far up the side of the store does the ladder reach? Answer: 11. 05 ft ***All I want to know is how to get to the answer, any help?

OpenStudy (anonymous):

1) Basically put the number of the ones you want over the whole. There are 10 blue marbles. There are 22 marbles total. Therefore, 10/22 = 5/11 2) Use trigonometry. sin 67 = x/12 12 sin 67 = x x ≈ 12(0.9205) x ≈ 11.05

OpenStudy (anonymous):

Thanks again. :)

OpenStudy (anonymous):